# Efficiency of Unsupervised Anomaly Detection Methods on Software Logs

### v1

@@ 3.1. AUC-ROC results
Words
IF
0.672
0.783 0.611
0.990
0.764
KMeans
0.722
0.293
0.579
0.971
0.641
RM
0.638
0.217
0.365
0.999
0.555
Trigrams
IF
0.675
0.814 0.698
0.981
0.792
KMeans
0.741
0.166
0.611
0.956
0.619
RM
0.735
0.102
0.540
0.999
0.594
Events
IF
0.755
0.693 0.889
0.978
0.829
KMeans
0.766
0.187
0.847
0.984
0.696
RM
0.581
0.222
0.486
0.929
0.555
Table 3. AUC with filtered (normal only) training data
@@
Words
OOVD
0.777
0.802
0.786
0.543
0.727
IF
0.482
0.433
0.611
0.954
0.620
KMeans
0.937
0.893
0.579
0.940
0.837
RM
0.958 0.967
0.381
0.950
0.814
Trigrams
OOVD
0.997 0.997 0.845
0.545
0.846
IF
0.605
0.530
0.845
0.944
0.731
KMeans
0.945
0.582
0.802
0.953
0.820
RM
0.988
0.714
0.619
0.982
0.826
Events
OOVD
0.995 0.895
0.903
0.535
0.832
IF
0.004
0.104
1.000
0.937
0.511
KMeans
0.041
0.235
0.806
0.981
0.515
RM
0.405
0.387
0.278
0.840
0.477
The results for AUC-ROC are showcased in Table 2 and Table 3. It is apparent that there is no single model-representation pair that gets the best results on each dataset. IF is a strong candidate for the best model if one only has unfiltered data contain both anomalies and normal observations without labels. KMeans performs similarly to IF on other datasets but fails completely on Thunderbird. RM performs poorly on the unfiltered setup as is to be expected due to the anomalies being in the vocabulary. The reason RM gets a good score with unfiltered HDFS is that the dataset just has little anomalies in general so when RM sees something rare it will get a higher anomaly score.
With the filtered (normal only) setup both KMeans and RM improve the results drastically when using the word- or trigram-representation. It is easy to intuitively understand why this happens on the RM. As the vocabulary only consists of what is normal, the model learns what is normal and deviations from that can be considered abnormal, not just rare. It is interesting that the performance went down with the filtered training data on both the IF model and the event-representation. Naturally, predicting based on the event ID introduces a lot of variation as BGL and Tb only have one item to predict with due to being event level labeled. Additionally, the modest preprocessing regexes might affect the parsing accuracy. Yet OOVD still managed to get good results with the events on BGL and Tb. This could suggest that the parser works well but other models are not able to handle the OOV events well.
The OOVD did well for everything except HDFS. This is likely due to the nature of the data. If we observe the values that OOVD and RM get, it is usually the case that when one does relatively well, the other does relatively worse. This is caused by the fact that RM (along with IF and Kmeans) do not have a way of distinguishing between OOV terms and other zeros in the matrices. In practice, the question that flips the balance is whether the term is never seen or very rarely seen. On BGL-trigrams both the RM and OOVD get great results which suggests that the anomalous log messages have both OOV and very rare terms in them. An example of this could be a folder path as the hierarchical structure goes from generic to rare by its nature. The folder example also explains why the OOVD did worse on the word-representation. As the words are not separated by slashes, the whole path becomes a long word that is likely to increase the anomaly score.
3.2. Time results
Table 4. Time with filtered training data (s)
@@ 3.2. Time results
Words
OOVD
1.29 5.30 ¡ 0.01
0.20
1.70
IF
34.00
185.87
0.13
9.49
57.37
KMeans
2.58
13.97
0.31
0.53
4.35
RM
1.62
9.27
¡ 0.01 0.19
2.77
Trigrams
OOVD
2.96
34.53
¡ 0.01 0.36
9.46
IF
93.24
999.70
0.15
48.46
285.39
KMeans
4.16
25.35
0.03
1.61
7.79
RM
2.02 11.66
0.01
0.43
3.53
Events
OOVD
0.74 4.31 ¡ 0.01 0.18 1.31
IF
21.90
106.43
0.17
4.53
33.26
KMeans
2.17
10.56
0.03
0.48
3.31
RM
1.37
7.83
0.01
0.23
2.36
There were no substantial differences between the runtimes on the unfiltered and filtered setups. Table 4 show that OOVD and RM are generally the two fastest approaches with KMeans being a little slower. Depending on the data size and use case, IF can become unfeasible due to the low speed. For example, with the trigram-representation IF could be 100 times slower than RM on certain datasets. Regarding log representation, the runtime is dependent on the number of items in the data. Therefore, trigrams are the slowest while events are the fastest.
If we account for pre-processing as shown in Table 5, it becomes apparent that the word-representation is the fastest end-to-end approach when starting with a raw log file. Generating trigrams turns out to be a slow process with the standard Python tools. In fact, it was slower than the parser. However, at the time of writing there is a pull request in Polars GitHub that will improve trigram generation likely making the speed closer to word creation speed.
Table 5. Pre-processing times (s)
@@
Load
2.77
7.86
24.71
372.24
Normalize
4.47
13.73
117.90
89.52
Create trigrams
9.50
142.58
378.16
861.33
Create words
0.38
1.83
9.03
15.80
Parse event IDs
4.69
137.18
273.67
716.53
4. Discussion
RQ1 regarded the AUC-ROC and time performance. As the results indicate, OOVD reached the best overall AUC-ROC score despite failing on HDFS. Timewise OOVD also performed the best with RM as a close second. In general the times seemed to better than suggested by previous work (Chen et al., 2021). For RQ2, we examined the log representation and found that trigram-representation can reach the best results but it is computationally much more expensive than the word-representation. Finally, for RQ3 we studied the impact of filtering out anomalies from the training data. It improved the overall results except for the IF model and the event-representation. In fact, IF model with events had the best average results for unfiltered training data and OOVD with trigrams for filtered data.
@@ 4. Discussion
Figure 1. Distribution of the RM scores on BGL (Word-representation, filtered setup)
@@
By using the AUC-ROC metric this study sheds light on methods that show potential but might have difficulty setting a threshold for anomaly detection. Figure 1 illustrates the distribution of RM scores for BGL using the word-representation. What the figure shows is that if we could assign a single value which would split the blue and red areas completely, we would have a perfect classifier. We can use such figures for aiding threshold selection. Of course in a real use case you would not see the colours.
@@
Consider visualizing both red and blue bars as black in Figure 1 and observe the significant gap between them, particularly the one occurring near 250. This gap represents an ideal point for setting a threshold. If we set the threshold at 250 as suggested by the figure we would get an accuracy of 0.957 and an F1-score of 0.721.
@@
This study opens various directions for future research. For example, finding a way to combine OOVD and with other models in a way that they do not clash with each others’ results has the promise to reach top scores on each dataset. The premise of this paper, fast and unsupervised anomaly detection, in itself encourages further studies to move toward more practical application of the anomaly detection approaches.
@@
As a limitation of the study, parameters are very crucial on unsupervised methods but systematic parameter optimization for the models was not in the scope of this paper. Additionally, the lack of F1-score as a metric means that the results are not directly comparable with similar previous studies (Chen et al., 2021), (Meng et al., 2019).
5. Conclusion
@@
This paper investigates speed and performance of anomaly detection using unlabeled data. Our findings indicate that our simple custom models (OOVD and Rarity Model) outperform established models (KMeans and IF) in terms of speed. In terms of performance (AUC-ROC), IF emerges as a viable option when dealing with unfiltered training data comprising both normal instances and anomalies. OOVD appears to be the preferred choice when the training data exclusively contains normal data. However, caution is advised since even the top-performing models exhibit subpar performance in certain scenarios. Nevertheless, these insights offer a good starting point for those aiming to conduct fast, unsupervised anomaly detection.

### v2

@@ 4-A AUC-ROC results
Words IF 0.672 0.783 0.611 0.990 0.764
KMeans 0.722 0.293 0.579 0.971 0.641
RM 0.638 0.217 0.365 0.999 0.555
Trigrams IF 0.675 0.814 0.698 0.981 0.792
KMeans 0.741 0.166 0.611 0.956 0.619
RM 0.735 0.102 0.540 0.999 0.594
Events IF 0.755 0.693 0.889 0.978 0.829
KMeans 0.766 0.187 0.847 0.984 0.696
RM 0.581 0.222 0.486 0.929 0.555
TABLE IV: AUC with Normal-only training data
@@
Words OOVD 0.777 0.802 0.786 0.543 0.727
IF 0.482 0.433 0.611 0.954 0.620
KMeans 0.937 0.893 0.579 0.940 0.837
RM 0.958 0.967 0.381 0.950 0.814
Trigrams OOVD 0.997 0.997 0.845 0.545 0.846
IF 0.605 0.530 0.845 0.944 0.731
KMeans 0.945 0.582 0.802 0.953 0.820
RM 0.988 0.714 0.619 0.982 0.826
Events OOVD 0.995 0.895 0.903 0.535 0.832
IF 0.004 0.104 1.000 0.937 0.511
KMeans 0.041 0.235 0.806 0.981 0.515
RM 0.405 0.387 0.278 0.840 0.477
Figure 1: The performance metrics on the Normal-only test setup on all datasets and representations
Figure 2: The performance metrics on the Unfiltered test setup on all datasets and representations
The results for AUC-ROC are showcased in Table III and Table IV. It is apparent that there is no single model-representation pair that gets the best results on each dataset. IF is a strong candidate for the best model for the Unfiltered scenario where data contains both anomalies and normal observations without labels. KMeans performs similarly to IF on other datasets but fails completely on Thunderbird. RM performs poorly on the Unfiltered data as is to be expected due to the anomalies being in the vocabulary. The reason RM gets a good score with Unfiltered HDFS is that the dataset just has few anomalies in general so when RM sees something rare it will get a higher anomaly score.
With the Normal-only scenario both KMeans and RM improve the results drastically when using the word- or trigram-representation. It is easy to intuitively understand why this happens on the RM. As the vocabulary only consists of what is normal, the model learns what is normal and deviations from that can be considered abnormal, not just rare. It is interesting that the performance went down with the Normal-only data on both the IF model and the event-representation. Naturally, predicting based on the event ID introduces a lot of variation as BGL and Tb only have one item to predict with due to being event level labeled. Additionally, the modest preprocessing regexes might affect the parsing accuracy. Yet OOVD still managed to get good results with the events on BGL and Tb. This could suggest that the parser works well but other models are not able to handle the OOV events quite well.
The OOVD did well for everything except HDFS. This is likely due to the nature of the data. If we observe the values that OOVD and RM get, it is usually the case that when one does relatively well, the other does relatively worse. This is caused by the fact that RM (along with IF and KMeans) do not have a way of distinguishing between OOV terms and other zeros in the matrices. Because the vectorizer is fitted on training data and transformed on test data, the terms that appear in the test data but not in the training data (OOV terms) are effectively ignored which is intended functionality for Sklearn vectorizers. In practice, the question that flips the balance is whether the term is never seen or very rarely seen.
On BGL-trigrams both the RM and OOVD get great results which suggests that the anomalous log messages have both OOV and very rare terms in them. An example of this could be a folder path as the hierarchical structure goes from generic to rare by its nature. The folder example also explains why the OOVD did worse on the word-representation. As the words are not split by slashes, the whole path becomes a long word that is likely to increase the anomaly score.
4-B F1-score results
The F1-score reached lower values than AUC-ROC despite using labels for the threshold selection. Figures 1 and 2 illustrate the results for each dataset and representation as bar plots for both F1-score and AUC-ROC. Fig. 1 visualizes well how, like with AUC-ROC, OOVD performs the best with Normal-only training data on F1-score as well. Similarly, Fig. 2 shows that IsolationForest remains the top model for the Unfiltered test setup.
Finally all the scores are summarized in Fig. 3 along with a fit line that describes the correlation between the AUC-ROC and F1 metrics. The Pearson correlation coefficient is 0.43 with p-value of $3.71~{}\times~{}10^{-5}$ which indicates statistically significant moderate correlation between the metrics. The correlation is apparent in cases where a model performed well on both metrics. There is also a cluster of cases where both performed poorly. The extent to which the F1-score and the AUC-ROC differ is largely due to the dataset. For example, in the plot the row of high F1-scores at the top are almost all from Hadoop. Similarly, the cluster of high AUC-ROC and low F1-scores are from Thunderbird and BGL.
Figure 3: All the results plotted by AUC-ROC and F1-score
4-C Time results
TABLE V: Time with filtered training data (seconds)
@@ 4-C Time results
Words OOVD 1.29 5.30 $<$0.01 0.20 1.70
IF 34.00 185.87 0.13 9.49 57.37
KMeans 2.58 13.97 0.31 0.53 4.35
RM 1.62 9.27 $<$0.01 0.19 2.77
Trigrams OOVD 2.96 34.53 $<$0.01 0.36 9.46
IF 93.24 999.70 0.15 48.46 285.39
KMeans 4.16 25.35 0.03 1.61 7.79
RM 2.02 11.66 0.01 0.43 3.53
Events OOVD 0.74 4.31 $<$0.01 0.18 1.31
IF 21.90 106.43 0.17 4.53 33.26
KMeans 2.17 10.56 0.03 0.48 3.31
RM 1.37 7.83 0.01 0.23 2.36
There were no substantial differences between the run times on the unfiltered and filtered setups. Table V show that OOVD and RM are generally the two fastest approaches with KMeans being a little slower. Depending on the data size and use case, IF can become infeasible due to the low speed. For example, with the trigram-representation IF could be 100 times slower than RM on certain datasets. Regarding log representation, the runtime is dependent on the number of items in the data. Therefore, trigrams are the slowest while events are the fastest.
If we account for pre-processing as shown in Table VI, it becomes apparent that the word-representation is the fastest end-to-end approach when starting with a raw log file. Generating trigrams turns out to be a slow process with the standard Python tools. In fact, it was slower than the parser. However, with further optimization, in Polars in particular, it should be possible to achieve speeds that are closer to word creation speed.
TABLE VI: Pre-processing times (seconds)
@@
Load 2.77 7.86 24.71 372.24
Normalize 4.47 13.73 117.90 89.52
Create trigrams 9.50 142.58 378.16 861.33
Create words 0.38 1.83 9.03 15.80
Parse event IDs 4.69 137.18 273.67 716.53
5 Discussion
RQ1 regarded the model performance in terms of AUC-ROC, F1-score and time. As the results indicate, OOVD reached the best overall AUC-ROC and F1-scores despite failing on HDFS. Time-wise OOVD also performed the best with RM as a close second. In general, the times were better than suggested by previous work [7]. For RQ2, we examined the log representation and found that trigram-representation can reach the best results but it is computationally much more expensive than the word-representation. Finally, for RQ3, we studied the impact of filtering out anomalies from the training data, Unfiltered vs. Normal-only scenario. Normal-only scenario improved the overall results except for the IF model and the event-representation. In fact, IF model with events had the best average results for Unfiltered training data and OOVD with trigrams for Normal-only data.
@@ 5 Discussion
Figure 4: Distribution of the RM scores on BGL (Word-representation, filtered setup)
@@
By emphasizing the AUC-ROC metric this study sheds light on methods that show potential but might have difficulty setting a threshold for anomaly detection. Fig. 4 illustrates the distribution of RM scores for BGL using the word-representation. What the figure shows is that if we could assign a single value which would split the blue and red areas completely, we would have a perfect classifier. Of course in a real use case you would not see the colours. We can use such figures for aiding threshold selection.
@@
Now, consider visualizing both red and blue bars as black in Fig. 4 and observe the surprising gap-up in the otherwise downward occurring near 250. This gap-up represents an ideal point for setting a threshold. If we set the threshold at 250 as suggested by the figure we would get an accuracy of 0.957 and an F1-score (binary) of 0.721. If we can automate this process, we could perform unsupervised threshold selection as well. However, due to the complexity of such process, it was out of scope of this paper, but it remains a possible topic for future research.
While the F1-score was considered as a secondary metric in this study, it provides a better connection to the existing literature such as Loglizer [7] and AutoLog [14] than only using AUC-ROC. Direct comparisons are still infeasible due to differences in preprocessing and label usage, but results show similarities, for example in the poor performance of the IF model on certain test cases. Further, this paper aims to contribute by presenting various models and representation methods for researchers that encounter low performance (as measured by F1-score) on dataset specific cases. Also, in metric selection itself, this study showcased that there is only a moderate correlation (0.43) between F1-score and AUCROC across all test cases. In essence, this means that generalizing the scores across datasets is difficult and calls for appropriate attention to metric selection.
While the Normal-only test scenario gained better results than the Unfiltered, it should be noted that there is a fundamental difference with these two approaches in what is actually being examined. In the Unfiltered approach, one is completely at the mercy of the data, and if there are a large number of anomalies, it is impossible to identify which are anomalies and which are not. The Unfiltered approach essentially relies on the assumption that rarity equates anomalousness. To some extent, the models can group similar rows under the same category which can cause a rare individual row to be correctly interpreted as normal, making the model potentially better than a simple frequency table with a threshold. Whereas, with the Normal-only data filtering, we do not need to equate rarity with anomalousness, but instead, the analysis is based on the deviation from normal. Then, the fundamental question is not what is rare enough but, rather, what is different enough. As such, we argue that in unsupervised learning, training with only normal data better suits anomaly detection while the unfiltered approach is closer to outlier detection. For academics, these results emphasize the importance of communicating the data selection process clearly.
@@
This study opens various directions for future research. For example, finding a way to combine OOVD with other models in a way that they do not clash with each others’ results has the promise to reach top scores on each dataset. Another direction to extend these results is to introduce new configurations for the data preprocessing, such as only using chronologically earlier data to predict later events as was done in a recent semi-supervised study [27]. The premise of this paper, fast and unsupervised anomaly detection, in itself encourages further studies to move toward more practical application of the anomaly detection approaches.
@@
As a limitation of the study, parameters are very crucial on unsupervised methods but systematic parameter optimization for the models was not in the scope of this paper. Additionally, utilizing the labels for the F1-score is not realistic and not necessarily comparable to those of similar previous studies [7], [12]. We found that the results were very stable on BGL, HDFS and Thunderbird, which warranted us to only report the results of just one run. For Hadoop, though, the results varied quite a bit. This is a limitation of the study that can be addressed in future research.
6 Conclusion
@@
This paper investigates speed and performance of anomaly detection using unlabeled data. Our findings indicate that our simple custom models (OOVD and Rarity Model) outperform established models (KMeans and IF) in terms of speed. In terms of performance (AUC-ROC, F1-score), IF emerges as a viable option when dealing with Unfiltered training data comprising both normal instances and anomalies. OOVD appears to be the preferred choice when the training data is Normal-only. However, caution is advised since even the top-performing models exhibit subpar performance in certain scenarios. Nevertheless, these insights offer a good starting point for those aiming to conduct fast, unsupervised anomaly detection.
Acknowledgment
Funded by the Academy of Finland (grant ID 349487).